Jasmine has 24 erasers and Muthu has 20 erasers.How many erasers must Muthu give to Jasmine so that she will have thrice the number of erasers that Muthu has?
1
Expert's answer
2011-10-13T12:41:11-0400
Let's mark number of Jasmine's erasers by J and the number of Muthu's erasers by M.
N - the unknown number of erasers.
J + N = 3(M - N)
J + N = 3M - 3N
4N = 3M - J
N = (3M - J)/4 = (3*20 - 24)/4 = 9.
So, Muthu must give 9 erasers to Jasmine.